### v1

# Quasi Inverse of Qubit Channels for Mixed States

Abstract
We found the quasi inverse of qubit channels as a unitary map, $\mathcal{E}^{i}$, by minimizing the average trace distance between the input state to the channel and the output of the quasi inverse channel for arbitrary qubit channel $\mathcal{E}$ and for arbitrary input states. The channel $\mathcal{E}$ was assumed completely positive and trace-preserving. To find the quasi inverse for mixed states, we proposed an alternative definition of the quasi inverse based on the mean square of the trace distance (MSTD) of the mixed input state of the original channel and the output of the quasi inverse. The definition based on the trace distance allowed easy generalization of the quasi inverse to mixed input states. The quasi inverse of the Pauli, generalized amplitude damping, mixed unitary, and tetrahedron channels calculated based on trace distance agreed with the one computed using fidelity in the special case of input states being pure.
preprint: APS/123-QED
I Introduction
Qubits are the building blocks of quantum computers, but they are quite vulnerable to unwanted interactions with environment. This interaction leads to the decoherence of the quantum state and makes the computations erraneous on quantum computers. This interaction process can be modeled as a quantum channel [1]. In general, quantum channels are irreversible operations [2, 3, 4]. However, the idea of quasi inversion of the quantum channel was recently introduced [2] to reverse a part of the effect of that channel. The concept of quasi inverse is essential for enhancing the fidelity of quantum gates and reducing errors in quantum computations. This idea is also helpful in quantum inverse problems, such as state estimation, where quasi inverses are used to improve efficiency and resilience [5].
The quasi inverse is defined so as to increase the average fidelity of the channel in Ref. [2]. The derivation of the quasi inverse was based on maximising the average fidelity between the input state of the channel and the output state of the quasi inverse of that channel. However, the derivation was restricted to the case when the input state was a pure state. In the present work, we propose an alternative way of defining the quasi inverse using trace distance that allows us to consider mixed input state.
The distance measures between the quantum states play a pivotal role in quantum information science. This is used to compute quantum correlations, entanglement, and coherence as discussed in [6]. The distance measures come into play for several critical tasks in quantum information processing. They help us gauge and distinguish different states of entanglement in bipartite and multipartite systems, especially when the observable set is not closed under products [7]. Among these measures, the trace distance and fidelity are frequently used [1, 8]. However, fidelity is usually easier to calculate only when one of the state is a pure state. On the other hand, the trace distance does not suffer from this drawback. Therefore, to extend the concept of quasi inverse for mixed input states, we propose the mean square trace distance (MSTD) averaged over all possible input states as a meteric to be minimized to find the inverse.
The structure of the paper is as follows: In Section II, we define the quasi inverse based on the mean of the trace distance and the general derivation of quasi inverse for qubit channels is presented in Section III. In Section IV, we derive the quasi inverses of a few qubit channels and concluding remarks are presented in Section V.
[…]
Figure 1: The schematic showing the qubit channel $\mathcal{E}$ with mixed input state $\rho$ and its quasi inverse $\mathcal{E}^{i}$.
II Definition of quasi inverse in terms of trace distance
Consider a general single-qubit channel $\mathcal{E}$ that converts a mixed state $\rho$ to $\zeta=\mathcal{E}(\rho)$. Suppose further that the quasi inverse of this channel is $\mathcal{E}^{i}$ that converts $\zeta$ to $\rho^{\prime}=\mathcal{E}^{i}(\zeta)$ as shown schematically in Fig. 1. The density matrices of the states $\rho$ and $\rho^{\prime}$ can be described in terms of the Bloch vectors $\mathbf{r}$ and $\mathbf{r}^{\prime}$, respectively, as $\rho=(I+\mathbf{r}\cdot\mathbf{\sigma})/2$ and $\rho^{\prime}=(I+\mathbf{r}^{\prime}\cdot\mathbf{\sigma})/2$ with $\mathbf{\sigma}$ being the vector of the Pauli operators [1]. Similarly, $\zeta=(I+\mathbf{z}\cdot\mathbf{\sigma})/2$.
We define the mean of the square of the trace distance (MSTD) between $\rho$ and $\rho^{\prime}$ as
$\overline{D^{2}}(\rho,\rho^{\prime})=\frac{1}{4}\int\absolutevalue{\mathbf{r}-\mathbf{r}^{\prime}}^{2}~{}\mathrm{d}^{3}\mathbf{r}\,,$ (1)
where the integral is taken over the unit Bloch ball with $r\leq 1$ to average over all mixed input states $\rho$. Let us note that this integral will be over the surface of the Bloch sphere if one wishes to consider only the pure input states $\rho=\ket{\psi}\bra{\psi}$ |. However, in this paper we assume that the input states can be mixed as well as pure.
The channel $\mathcal{E}^{i}$ will be quasi inverse if it minimizes MSTD in Eq. (1) in the optimal way, regardless of the input state. Therefore, similar to the definition of the quasi inverse based on fidelity [2], we define the quasi inverse $\mathcal{E}^{i}$ as a channel that reduces the MSTD between the state $\rho$ and the output state of the channel by at least the same or more than any other channel $\mathcal{E}^{\prime}$, i.e.,
$\overline{D^{2}}(\mathcal{E}^{i}(\zeta),\rho)\leq\overline{D^{2}}(\mathcal{E}^{\prime}(\zeta),\rho)\qquad\forall\quad\mathcal{E}^{\prime}$ (2)
It was proved in Ref. [2] that the inverse channel can only be a unitary operator if the input state is assumed to be a pure state and the inverse is defined via average fidelity. Let us note that this can be assumed to be true in general as well about the quasi inverse as follows: The action of a quantum channel can be decomposed into rotation and scaling of the Bloch vector of the input state [1]. Since any trace-preserving quantum channel cannot increase the length of the Bloch vector, the optimal quantum operation to reverse the effect of rotation and scaling can only be rotation embodied by a unitary transformation. Therefore, the inverse channel essentially rotates the state towards the input state to reduce its distance from the original state without changing the length of the Bloch vector. So, we can assume that the optimal quasi inverse can only be a unitary operation.
Assuming the quasi inverse as a unitary operator, we can write
$\mathcal{E}^{i}(\zeta)=V\zeta V^{\dagger}\,,$ (3)
where $V$ is a unitary operator that can be written as
$V=x_{0}\mathbb{I}+i\mathbf{x}\cdot\mathbf{\sigma}$ (4)
with $x_{0}$ a real number, $\mathbf{x}$ real vector, and $x_{0}^{2}+\mathbf{x}\cdot\mathbf{x}=1$.
Finding the quasi inverse is essentially finding four real paramters $(x_{0},\mathbf{x})$ that specifies the unitary $V$. To do this, we can setup a constrained optimization problem to maximise $\Delta\overline{D^{2}}$ (the decrease in the MSTD)
$\Delta\overline{D^{2}}=\overline{D^{2}}(\rho,\zeta)-\overline{D^{2}}(\rho,\rho^{\prime})\,$ (5)
over all unitary maps, i.e., maximizing over the real parameters $(x_{0},\mathbf{x})$, subject to the constraint $x_{0}^{2}+\mathbf{x}\cdot\mathbf{x}=1$, for a given channel $\mathcal{E}$.
III General Derivation
When the state passes through an arbitrary channel $\mathcal{E}$, the Bloch vector of the qubit state is transformed as [1]
$\mathbf{r}\longrightarrow\mathbf{z}=M\mathbf{r}+\mathbf{c}\,,$ (6)
where $\mathbf{M}$ is a real $3\times 3$ matrix and $\mathbf{c}$ is a real vector in $\mathbb{R}^{3}$ given by
$M_{ij}=\frac{1}{2}\text{Tr}(\sigma_{i}\mathcal{E}(\sigma_{j}))\,,\qquad c_{i}=\frac{1}{2}\text{Tr}(\sigma_{i}\mathcal{E}(\mathbb{I}))\,.$ (7)
Hence, a quantum channel can be characterized by the pair $(\mathbf{M},\mathbf{c})$, and composition of quantum channel is reflected in the composition of the transformation matrices [1]
$\mathcal{E}_{2}\circ\mathcal{E}_{1}\equiv(M_{2}M_{1},M_{2}\mathbf{c}_{1}+\mathbf{c}_{2})$ (8)
with $\mathcal{E}_{1,2}\equiv(M_{1,2},\mathbf{c}_{1,2})$.
Using Eq. (1), the MSTD of the channel between $\rho$ and $\zeta$ is found as
$\overline{D^{2}}(\rho,\zeta)=\frac{1}{20}\left(\text{Tr}(MM^{\dagger})-2\text{Tr}(M)+3\right)+\frac{1}{4}\absolutevalue{\mathbf{c}}^{2}\,.$ (9)
To evaluate the MSTD integral, the following identities were used:
$\int r_{i}~{}\mathrm{d}V=0\,,\qquad\int r_{i}r_{j}~{}\mathrm{d}V=\frac{1}{5}\delta_{ij}\,,$ (10)
$\begin{gathered}\mathrm{d}V=r^{2}\sin\theta~{}\mathrm{d}r~{}\mathrm{d}\theta~{}\mathrm{d}\phi\,,\\ \mathbf{r}=\langle r\sin\theta\cos\phi,r\sin\theta\sin\phi,r\cos\theta\rangle\,,\\ r\in[0,1]\,,\quad\theta\in[0,\pi]\,,\quad\phi\in[0,2\pi]\,.\end{gathered}$ (11)
The MSTD of combined channel $\mathcal{E}^{i}\circ\mathcal{E}$ is given by
$\overline{D^{2}}(\rho,\rho^{\prime})=\frac{1}{20}\left(\text{Tr}(NN^{\dagger})-2\text{Tr}(N)+3\right)+\frac{1}{4}\absolutevalue{\mathbf{u}}^{2}\,,$ (12)
where $N=M^{i}M$ and $\mathbf{u}=M^{i}\mathbf{c}+\mathbf{c}^{i}$.
Substituting Eqs. (9) and (12) in Eq. (5) and maximization process gives out the quasi inverse for any qubit channel with mixed input states.
IV Examples
To find the qusi inverse of specific channels, we consider the same example channels as in Ref. [2] for comparison of the results.
IV.1 The Pauli Channel
Pauli channel is ubiquitous in quantum information processing devices and is used as the most common model for qubit noise in computing hardware. It can be specified using Kraus operators as
$\mathcal{E}(\rho)=p_{0}\rho+\sum_{i=1}^{3}p_{i}\sigma_{i}\rho\sigma_{i}$ (13)
with $p_{i}\geq 0$ and $\sum_{i=0}^{3}p_{i}=1$. This leads to
$\Delta\overline{D^{2}}=\frac{2}{5}\begin{pmatrix}x_{0}&\mathbf{x}^{T}\end{pmatrix}Q\begin{pmatrix}x_{0}\\ \mathbf{x}\end{pmatrix}\,,$ (14)
where
$\begin{gathered}Q=\text{diag}\left(0,p_{1}-p_{0},p_{2}-p_{0},p_{3}-p_{0}\right)\,.\end{gathered}$ (15)
The maximum value of decrease in MSTD is given by
$\Delta\overline{D^{2}}=\frac{2}{5}\max(\lambda_{\max},0)\,,$ (16)
where $\lambda_{\max}$ is the largest eigenvalue of matrix $Q$ in Eq. (15). The normalized eigenstate $(x_{0},~{}\mathbf{x})^{T}$ corresponding to this largest eigenvalue will determine the quasi inverse which is given by
$V=i\sigma_{i}\quad\text{ if ~{}}p_{i}=p_{\max}\quad i\in\{1,2,3\}\,.$ (17)
In this case, the decrease in the MSTD is given by
$\Delta\overline{D^{2}}=\frac{2}{5}\max(p_{\max}-p_{0},0)\,,$ (18)
where $p_{\max}=\max(p_{1},p_{2},p_{3})$. Therefore, if $p_{\max}>p_{0}$, the quasi inverse of the Pauli channel exists. We find that if $p_{0}\leq 1/2$ and $p_{\max}\geq 1/2$, then $\Delta\overline{D^{2}}$ is maximized.
IV.2 Generalized Amplitude Damping Channel
The generalized amplitude damping channel $\mathcal{E}_{\text{GAD}}$ is a non-unital channel characterized by Krauss operators [1]
$E_{1}$ $=\sqrt{p}\begin{pmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{pmatrix}\,,$ (19)
$E_{2}$ $=\sqrt{p}\begin{pmatrix}0&\sqrt{\gamma}\\ 0&0\end{pmatrix}\,,$
$E_{3}$ $=\sqrt{1-p}\begin{pmatrix}\sqrt{1-\gamma}&0\\ 0&1\end{pmatrix}\,,$
$E_{4}$ $=\sqrt{1-p}\begin{pmatrix}0&0\\ \sqrt{\gamma}&0\end{pmatrix}\,.$
This leads to
$\Delta\overline{D^{2}}=\frac{2}{5}\begin{pmatrix}x_{0}&\mathbf{x}^{T}\end{pmatrix}Q\begin{pmatrix}x_{0}\\ \mathbf{x}\end{pmatrix}\,,$ (20)
where
$Q=\frac{1}{2}\text{diag}\big{(}0,-\alpha(\alpha+1),-\alpha(\alpha+1),-2\alpha\big{)}$ (21)
and $\alpha^{2}={1-\gamma}$.
Clearly, the largest eigenvalue $\lambda_{\max}=-\alpha$ for $\alpha<0$ and thus the quasi inverse is
$V=i\sigma_{3}\,.$ (22)
In this case, the maximum decrease in MSTD is given by
$\Delta\overline{D^{2}}=\frac{2}{5}\begin{cases}0\,,&\quad\text{for }\alpha>0\,,\\ -\alpha\,,&\quad\text{for }\alpha<0\,.\end{cases}$ (23)
IV.3 Mixed Unitary Channel
A mixed unitary channel is given by [2]
$\mathcal{E}(\rho)=(1-3p)\rho+p\sum_{i}^{3}U_{i}\rho U_{i}^{\dagger}\,,$ (24)
where $U_{i}=\exp(-i\frac{\theta}{2}\sigma_{i})=\cos(\frac{\theta}{2})I-\sin(\frac{\theta}{2}\sigma_{i})$ is a unitary operator that rotates the qubit around $x_{i}$ axis by an angle $\theta$. This leads to
$\Delta\overline{D^{2}}=\frac{2}{5}\begin{pmatrix}x_{0}&\mathbf{x}^{T}\end{pmatrix}Q\begin{pmatrix}x_{0}\\ \mathbf{x}\end{pmatrix}\,,$ (25)
where
$Q=\begin{pmatrix}0&v/2&v/2&v/2\\ v/2&q&0&0\\ v/2&0&q&0\\ v/2&0&0&q\end{pmatrix}\,,$ (26)
$v=p\sin(\theta)$, and $q=4p\sin[2](\frac{\theta}{2})-1$. For $q\geqslant 0$, the largest eigenvalue of this matrix is $\lambda_{\max}=\frac{1}{2}(q+$ $\sqrt{q^{2}+3v^{2}}$ ), with the corresponding eigenvector given by $\left(\frac{3v}{2\lambda_{\max}}\quad 1\quad 1\quad 1\right)^{T}$. This means that the quasi inverse of the channel is
$V=e^{i\phi\mathbf{n}\cdot\mathbf{\sigma}}\,,$ (27)
where
$\cos\phi=\frac{\sqrt{3}v}{\sqrt{3v^{2}+4\lambda_{\max}^{2}}},\quad\mathbf{n}=\frac{1}{\sqrt{3}}(x+y+z)\,.$ (28)
The decrease in MSTD is given by
$\Delta\overline{D^{2}}=\frac{2}{5}\lambda_{\max}\,.$ (29)
IV.4 Tetrahedron Channel
This is the channel whose quasi inverse is different from one of its own Kraus operators. Its mathematical form is given as
$\mathcal{E}(\rho)=q\rho+\sum_{i=0}^{3}p_{i}(\mathbf{v}_{i}\cdot\mathbf{\sigma})\rho(\mathbf{v}_{i}\cdot\mathbf{\sigma})\,,$ (30)
where
$q=1-\sum_{i=0}^{3}p_{i}$
and the vectors $\mathbf{v}_{i}$ are the corners of tetrahedron:
$\mathbf{v}_{0}$ $=\frac{1}{\sqrt{3}}\begin{pmatrix}1&1&1\end{pmatrix}\,,$ (31)
$\mathbf{v}_{1}$ $=\frac{1}{\sqrt{3}}\begin{pmatrix}1&-1&-1\end{pmatrix}\,,$
$\mathbf{v}_{2}$ $=\frac{1}{\sqrt{3}}\begin{pmatrix}-1&1&-1\end{pmatrix}\,,$
$\mathbf{v}_{3}$ $=\frac{1}{\sqrt{3}}\begin{pmatrix}-1&-1&1\end{pmatrix}\,.$
We consider a special case [2] with
$p_{1}=p_{2}=p\,,\quad p_{0}=p_{3}=p^{\prime}\,,$ (32)
where $p+p^{\prime}\leq 0.5$ due to normalization of probability. This leads to:
$\Delta\overline{D^{2}}=\frac{2}{5}\begin{pmatrix}x_{0}&\mathbf{x}^{T}\end{pmatrix}Q\begin{pmatrix}x_{0}\\ \mathbf{x}\end{pmatrix}\,,$ (33)
where
$Q=\begin{pmatrix}0&0&0&0\\ 0&\frac{8p}{3}+\frac{8p^{\prime}}{3}-1&-\frac{2p}{3}+\frac{2p^{\prime}}{3}&0\\ 0&-\frac{2p}{3}+\frac{2p^{\prime}}{3}&\frac{8p}{3}+\frac{8s}{3}-1&0\\ 0&0&0&\frac{8p}{3}+\frac{8p^{\prime}}{3}-1\end{pmatrix}\,.$ (34)
The maximum value of decrease in MSTD is given by
$\Delta\overline{D^{2}}=\frac{2}{5}\begin{cases}\displaystyle{\max}\left\{2p^{\prime}-1+\frac{10p}{3},0\right\}\,,&\text{ if }p\geq p^{\prime}\,,\\ \displaystyle{\max}\left\{2p-1+\frac{10p^{\prime}}{3},0\right\}\,,&\text{ if }p\leq p^{\prime}\,.\end{cases}$ (35)
The quasi inverse for such channel is thus given by:
$V=\begin{cases}\displaystyle\frac{\sigma_{1}+\sigma_{2}}{\sqrt{2}}\,,&\text{ if }\lambda_{\max}=2p^{\prime}-1+\displaystyle\frac{10p}{3}\,,\\[14.22636pt] \displaystyle\frac{\sigma_{1}-\sigma_{2}}{\sqrt{2}}\,,&\text{ if }\lambda_{\max}=2p-1+\displaystyle\frac{10p^{\prime}}{3}\,.\end{cases}$ (36)
Let us note that the quasi inverses of the four example channels in Eqs. (17), (22), (27), and (36) have turned out to be the same as in Ref. [2] using the fidelity. However, the expressions in Eqs. (18), (23), (29), and (35) for the decrease in the MSTD are different than those for pure states in Ref. [2] but reduce to the same value if we had performed averaging just over the surface of the Bloch sphere.
V Conclusions
We proposed an alternative definition for quasi inverse of a channel that can easily extend the quasi inverse of channels to mixed input states. The quasi inverse based on maximizing the mean squared trace distance (MSTD) over the full Bloch ball of input state of a completely positivie and trace preserving channel was postulated to be a unitary operator. This resulted in a constrained optimization program to find the four real paramters of the inverse operator. The results for Pauli, generalized amplitude damping, mixed unitary, and the tetrahadron channels agreed with those derived based on the fidelity maximization of a channel. Therefore, this new definition provides an equivalent methodology to find the quasi inverse that is also generalizable to mixed state.